# Balancing Confidentiality and Transparency for Blockchain-based Process-Aware Information Systems

### v2

@@ Abstract
Blockchain enables novel, trustworthy Process-Aware Information Systems (PAISs) by enforcing the security, robustness, and traceability of operations. In particular, transparency ensures that all information exchanges are openly accessible, fostering trust within the system. Although this is a desirable property to enable notarization and auditing activities, it also represents a limitation for such cases where confidentiality is a requirement since interactions involve sensible data. Current solutions rely on obfuscation techniques or private infrastructures, hindering the enforcing capabilities of smart contracts and the public verifiability of transactions. Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs aimed at preserving both confidentiality and transparency. Smart contracts enact, enforce and store public interactions, while attribute-based encryption techniques are adopted to specify access grants to confidential information. We assess the security of our solution through a systematic threat model analysis and assess its practical feasibility by gauging the performance of our implemented prototype in different scenarios from the literature.
@@
Blockchain enables new forms of trustable Process-Aware Information Systems (PAISs) and inter-organizational collaborations [36]. This is possible thanks to the characteristics of public permissionless blockchains, which provide strong guarantees removing the need for third-party authorities [30, 31]. Starting from the foundation, the distributed ledger and consensus mechanisms permit distributed and immutable information storage. These, combined with the cryptographic primitives, ensure security properties and accountability of recorded operations. Smart contracts are programs immutably stored inside the blockchain that can encode and enforce business logic. On top of this, transparency is a key aspect, as information stored in the ledger and performed operations are accessible to everyone within the system.
Business logic enforcement and transparency are highly desirable properties for collaborative process execution and monitoring [13, 14]. However, public observability of the whole set of exchanged data hinders the adoption of blockchain-based solutions in contexts where confidentiality is a critical requirement. Striking a balance between these two aspects is crucial in highly regulated sectors wherein sensitive data are treated (consider, for instance, the domains of pharmaceutical supply-chain and healthcare [18, 2]). Existing solutions to solve this conundrum employ private permissioned blockchains [39, 7]. However, these solutions may hide data from auditors and assume high robustness and security guarantees of the consortium’s system. Other approaches propose encryption techniques to restrict visibility of data stored on-chain only to authorized parties [28, 29, 23, 24]. However, their integration with process management systems is limited or non-existing.
Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs that preserves confidentiality of exchanged information while keeping public enforcement and transparency of process execution. To this end, we bridge the silos of blockchain technologies and encryption with business process management. In particular, we rely on smart contracts to encode and execute business process logic while logging the interactions between parties, and we resort to Attribute-Based Encryption (ABE) [34] to control the access of different parties to the activities’ data payloads and information artifacts, thus safeguarding confidentiality. We demonstrate the security of our approach against a threat model of the proposed architecture. Following the principles of Design Science Research (DSR), we implemented our artifact and evaluated the feasibility of our solution by showing its application in the context of a healthcare scenario taken from the literature and analyzing the execution performance with collaborative processes presented in the related scientific literature.
@@
2 Example and Problem Illustration
To illustrate the problem we aim to tackle, we introduce a running example in the healthcare domain inspired by [9]. The model depicted in Fig. 1 depicts the management process for X-ray diagnostics in the form of a BPMN choreography diagram [11]. The patient first presents the medical prescription, asking for an appointment. The radiology department checks the availability of the ward and, if a date is available, the appointment is confirmed. Otherwise, there is a new scheduling tentative. After obtaining the registration details, the temperature and vaccine certification of the patient are verified. If the temperature is under the threshold and the certification is valid, a radiology clerk checks the appointment identifier of the patient and returns the confirmation. If the appointment is confirmed, the X-ray exam is done and the results are provided to the ward with a final report. Finally, the patient receives the results from the ward.
As the example shows, ensuring accurate tracking of medical procedures is critical to improving process efficiency, trust in the public towards the healthcare system, and enabling auditing in case of misconduct or compliance with regulations. Data-based evidence is crucial to this end. The immutability, accessibility and persistence guarantees offered by a public-blockchain-based PAIS are fit for purpose from this standpoint. However, sensitive patient data must also be kept shielded from unauthorized inspections: publicly disclosing the result of the diagnostic analysis or information about vaccine certification would severely violate the privacy sphere of the patient, for example. Balancing a secure, publicly verifiable and decentralized infrastructure with the need for data confidentiality remains therefore an open challenge, which motivates our research.
@@ 3 Background and State of the Art
In this section, we first describe the main concepts on which our approach relies, namely blockchain technology and Multi-Authority Attribute-Based Encryption (MA-ABE), and then illustrate how those pillars constitute the foundations of existing approaches in the literature that are of relevance to our investigation.
@@ 3.1 Background
A ledger is an append-only singly-linked list of transactions, namely the transfer of assets among accounts. Every account is linked to a pair of keys: a public one for recognition and addressing, and a public one for the unlocking and expenditure of the account’s balance. A blockchain collates consecutive segments of the ledger in its body, while the header includes, among others, the number of the current block and the hash of the current and the previous block. The hash-link to the previous block preserves the total order of transactions in the ledger [12]. This hash cryptographically ensures that previous blocks, and the data therein, are not changed since this would corrupt the hash-links. Transactions can thus be included in an append-only manner, and it is not possible to delete old interactions. New blocks are added to the chain at regular intervals of time by selected computational nodes within the blockchain infrastructural network. Those nodes are needed to compute and finalize the blocks’ hash codes, among other things. New cryptocurrency is minted and distributed to the nodes’ accounts to pay back their computational effort. The entire blockchain data structure is stored and replicated among all nodes in the network, who have full visibility over the saved ledger. The use of consensus protocols and cryptographic primitives permits the convergence and public verification of the transactions included in every block of each replica, thereby guaranteeing decentralization and trustworthiness without the need for an independent third-party authority certification or governance. Several platforms implement the blockchain protocol (e.g., Ethereum [41]). In the last few years, blockchain technologies have been adopted in many different contexts, especially after the introduction of smart contracts, namely programs deployed onto the blockchain and executed within a dedicated virtual machine (e.g., the Ethereum Virtual Machine, EVM). Their functions are invoked through transactions. Core features of smart contracts are the immutability of their code and the storage of their invocations on the blockchain, making such information available for auditing purposes [15]. A smart contract has an account in which it can store and update data alongside its cryptocurrency balance. Smart contract executions incur a cost on the invoker. The cost is measured in gas units and depends on the complexity of the called function and the size of the exchanged data. To save on the latter, distributed tamper-proof storage systems are often employed to store the actual data in a non-modifiable file, while a permalink to that file is saved in the smart contract state to the datastore’s locator. The InterPlanetary File System (IPFS) [3] is a typical example of such a system. Based on a peer-to-peer network wherein each node stores chunks of the data, IPFS binds every file with a string used both as a unique identifier and resource locator. The string is based on the hash of the file content itself so that if the file is altered, the permalink does not match. The code execution is validated in each node of the blockchain, providing verifiable proof of the obtained results for the final users [20]. As we will see next, these features fostered the development of blockchain-based PAISs using smart contracts as the backbone for business (process) rule enforcing (in our example, the BPMN choreography is translated into smart contract code, and its steps correspond to invocations to its functions) [9, 13].
@@
Attribute-Based Encryption (ABE) is a type of public key encryption where the ciphertext, namely an encrypted plaintext, and the corresponding decryption key are linked through custom boolean attributes [34]. Ciphertext-Policy Attribute-Based Encryption (CP-ABE) [5], a variant of ABE, prescribes that every user is endowed with a set of attributes encoded in their key. For example, user 0xB0…1AA1 can be endowed with attributes Patient to specify their role and attribute PID476948 to indicate they are involved in process instance .
Data owners write policies to be attached to the CP-ABE-encrypted document (the ciphertext). Those policies express conditions for granting in-clear access to the data. Policies are propositional logic formulae employing ABE attributes as literals. The policy is evaluated on the attributes associated with a potential reader: Only if the formula is satisfied, the user’s key can decrypt the data.
@@
Attributes are typically assigned by one authority:
@@ 3.2 State of the Art
Over the last few years, several solutions that automate collaborative processes using blockchain technology have been designed [13, 8]. Several works in this field have demonstrated the effectiveness of blockchain-based solutions in improving trust among participants in multi-party collaborations [40] even in adversarial settings [26], enhance verifiability of workflows with model-driven approaches [25, 38], allow for monitoring [15], mining [21], security [23], and auditing [10, 7]. These studies enhance the integration of blockchain technology with process management, unlocking security and traceability benefits. However, they primarily focus on the control-flow perspective and lack mechanisms for secure access control to the stored data on public platforms.
A relevant area of research involves allowing for data privacy and integrity within blockchain systems. Several studies examine the application of encryption techniques to achieve this goal. Hawk [22] employs user-defined private Smart Contracts to automate the implementation of cryptographic protocols. Abid et al. [1] leverage Zero-Knowledge Proof (ZKP) and Fully Homomorphic Encryption (FHE) to build a solution that supports privacy-preserving traceability and the analysis of encrypted data. They store on-chain only the hash of the data used to verify the correctness, while all the computations and the data are stored off-chain. Benhamouda et al. [4] use a public blockchain storage layer for confidential data. Peng et al. [32] propose a consortium blockchain as a decentralized storage system for file sharing among organizations. Wu et al. [42] present an attribute-based access control scheme that preserves the privacy of both attributes and policies. Pham et al. [33] propose a decentralized storage solution that integrates InterPlanetary File System (IPFS), MA-ABE, and blockchain technology. Hong et al. [19] introduce a data-sharing solution based on decentralized ABE (MA-ABE), blockchain, and IPFS. Bramm et al. [6] present BDABE, an access control mechanism that applies a distributed ABE scheme within a consensus-driven infrastructure for real-world applications. Feng et al. [16] combine attribute-based encryption ABE and identity-based encryption (IBE) to achieve efficient data sharing and verification of data correctness. They employ blockchain technology to ensure tamper-proof and regulated data storage. Yan et al. [43] introduce a scheme for fine-grained access control by implementing proxy encryption and decryption while supporting policy hiding and attribute revocation. The encrypted data is stored on IPFS, and the metadata is stored on the blockchain. All these studies, along with the aforementioned [29, 28, 27], allow for secure access control on blockchains but lack the integration of this technology with process management systems.
@@ 4 The CONFETTY architecture
In this section, we describe our proposed architecture, named CONFidentiality EnforcemenT TransparencY (CONFETTY). We first introduce its key functionalities and the core components that support them (Sect. 4.1), then describe how these components interact and manage information to support confidential-preserving blockchain-based process execution (Sect. 4.2). Due to space restrictions, we do not enter the technical details of our approach but provide an overarching presentation of our solution’s rationale and key concepts. Finally, we conclude the section with some considerations on the benefits brought by our approach merging blockchain technologies and ABE.
@@ 4.1 Functional viewpoint
Figure 2 outlines the CONFETTY architecture,
@@
(i) The configure functionality foresees the specification of the process by the process owner. It requires the registration of process information including activities to execute, data to exchange, control-flow routing, the expected participants’ roles (e.g., the information enclosed in the choreography diagram in Fig. 1) and how they can operate with confidential data (via ABE policies) to control information access.
(ii) The instantiate functionality lets certifiers authenticate the participants, define their role in the process (e.g., that user 0xB0…1AA1 is a patient), and also kickstart the process run on-chain, by setting its public state.
(iii) The transact functionality allows process participants to interact following the process to exchange data at run time. This functionality manages both public and confidential data, updating and enforcing the process state according to the logic in the blockchain.
@@
The functionalities above are realized by four main components of CONFETTY: (i) The Process Manager handles all the operations involving the participants and related to the process. This is done by acting as an interface for the blockchain, abstracting from the technical implementation and exposing functionalities to the users. In particular, the Process Manager instantiates the process on the blockchain on given process specification data. To support the instantiate functionality, it provides capabilities for role assignment to participants. (ii) The Process Contract is a smart contract that manages the process instances and their execution. It elaborates and manages all process specifications and their instances’ public states, acting as a software realization both of factory and proxy patterns [17] akin to [10]. It supports the instantiate functionality by enforcing the control flow, the data flow, the assignment of tasks to participants, and the routing of decision points. In this way, the Process Contract updates the public state of the process and generates transparent records that are accessible and verifiable by interested auditors. (iii) The Confidentiality Manager is in charge of handling the participants’ authorizations and the operations over confidential data. It operates as a facade towards the blockchain to store access grants, distributes decryption keys to the users solely based on their roles and involvement in the process, and handles the storage and retrieval of encrypted data. (iv) The Confidentiality Contract is a smart contract handling the notarization of authorizations for confidential data, as well as for their writing and access. In addition to the above inner components, the architecture relies on two external ones: a Programmable Blockchain Platform (like Ethereum) and a Tamper-proof Distributed File Storage (like IPFS). The former maintains the public state, allowing for the creation of logic and constraints while notarizing the executed operations. It comprises two core elements: The Distributed Ledger, storing the transactions, and the Virtual Machine, hosting and running smart contracts. The latter (which we will henceforth name as Distributed FS for brevity), is used to store business data and thus improve the performance and costs of a blockchain. The data is saved on the external storage (upon encryption, if it must remain secret), while a resource locator with the hash of that data is kept on-chain for notarization and future retrieval.
@@
We conclude this subsection with a few considerations about the actors and trust model. We assume that the aforementioned users (process owners, certifiers, process participants, and auditors) possess a blockchain account with a key pair to be able to publicly identify themselves (as with the aforementioned user 0xB0…1AA1) and sign blockchain transactions built through the CONFETTY modules. We also remark that we consider the CONFETTY platform as a trusted party, while users are assumed to be honest but curious. Finally, notice that the involvement of a pre-appointed third-party certifier to authenticate process participants is necessary as a user alone can confirm their identity (e.g., via the aforementioned public-private key scheme) but cannot attest alone to the truthfulness of assertions regarding themselves.

### v3

@@ Abstract
Blockchain enables novel, trustworthy Process-Aware Information Systems (PAISs) by enforcing the security, robustness, and traceability of operations. In particular, transparency ensures that all information exchanges are openly accessible, fostering trust within the system. Although this is a desirable property to enable notarization and auditing activities, it also represents a limitation for such cases where confidentiality is a requirement since interactions involve sensitive data. Current solutions rely on obfuscation techniques or private infrastructures, hindering the enforcement capabilities of smart contracts and the public verifiability of transactions. Against this background, we propose CONFETTY, an architecture for blockchain-based PAISs to preserve confidentiality and transparency. Smart contracts enact, enforce and store public interactions, while attribute-based encryption techniques are adopted to specify access grants to confidential information. We assess the security of our solution through a systematic threat model analysis and evaluate its practical feasibility by gauging the performance of our implemented prototype in different scenarios from the literature.
@@
Blockchain enables new forms of trustable Process-Aware Information Systems (PAISs) and inter-organizational collaborations [27]. This is possible thanks to the characteristics of public permissionless blockchains, which provide strong guarantees removing the need for third-party authorities [22]. Starting from the foundation, the distributed ledger and consensus mechanisms permit distributed and immutable information storage. These features, combined with the cryptographic primitives, ensure security properties and accountability of recorded operations. Smart contracts are programs immutably stored inside the blockchain that can encode and enforce business logic. On top of this, transparency is a key aspect, as information stored in the ledger and performed operations are accessible to everyone within the system.
Business logic enforcement and transparency are highly desirable properties for collaborative process execution and monitoring [10]. However, public observability of the whole set of exchanged data hinders the adoption of blockchain-based solutions in contexts where confidentiality is a critical requirement. Striking a balance between these two aspects is crucial in highly regulated sectors wherein sensitive data are treated (consider, for instance, the domains of the pharmaceutical supply-chain and healthcare [11]). Existing solutions to solve this conundrum employ private permissioned blockchains [30, 6]. Nevertheless, these solutions may hide data from auditors and assume high robustness and security guarantees of the consortium’s system. Other approaches propose encryption techniques to restrict visibility of data stored on-chain only to authorized parties [15, 16, 20, 21]. However, their integration with process management systems is limited or non-existent.
Against this background, we propose an approach and an architecture for blockchain-based process enactment that preserves the confidentiality of exchanged information while keeping public enforcement and transparency of process execution. We name our approach CONFidentiality EnforcemenT TransparencY (CONFETTY). In particular, we rely on smart contracts to encode and execute business process logic while logging the interactions between parties, and we resort to Multi-Authority Attribute-Based Encryption (MA-ABE) [4] to control the access of different parties to the activities’ data payloads and information artifacts, thus safeguarding confidentiality. We demonstrate the security of our approach against a threat model for the proposed architecture. We implemented our artifact and evaluated the feasibility of our solution by showing its application in the context of a healthcare scenario taken from the literature. Also, we analyzed the execution performance with collaborative processes presented in the related scientific literature.
@@
2 Example, Problem Illustration and Requirements
To illustrate the problem we aim to tackle, we introduce a running example in the healthcare domain inspired by [7]. Figure 1 depicts the management process for X-ray diagnostics in the form of a BPMN choreography diagram. The patient first presents the medical prescription, asking for an appointment. The radiology department checks the availability of the ward, and if a date is available, the appointment is confirmed. Otherwise, there is a new tentative. After obtaining the registration details, the temperature and vaccine certification of the patient are verified. A radiology clerk checks whether the appointment can be confirmed. If so, the X-ray exam is done, and the results are provided to the ward with a final report. In the end, an insurance agency receives the patient’s results. The insurance agency can access the patient’s medical report to follow up with compensatory actions in the subscribed contract (e.g., the reimbursement of medical bills). Beyond the boundaries of the choreography diagram in Fig. 1, we assume the presence of an inspector of the local Ministry of Health. They are not among the process participants, yet they should have full access to the information exchanged even after the conclusion of the instance for auditing purposes.
Table 1 lists the requirements derived from the above use case. To begin with, ensuring public execution and tracking of healthcare procedures is critical to improving process efficiency and trust in the public towards the healthcare system. We concretize this aspect in Table 1. Enabling auditing of those procedures is crucial since it permits investigations on misconduct or lack of compliance with regulations (Table 1). This is allowed by the data-based evidence, which should be ensured by information storage in permanent, tamper-proof and non-repudiable information artifacts (Table 1). The guarantees offered by a public blockchain-based PAIS are fit for purpose from this standpoint. However, sensitive patient data must be shielded from unauthorized inspections (Table 1). For instance, although the insurance agency is involved in the process, it should not have full access to all exchanged data (e.g., the vaccine certification). Publicly disclosing the result of the diagnostic analysis or information about vaccine certification would severely violate the privacy sphere of the patient. To prevent this, access should be restricted only to designated and authorized entities through a fine-grained access control mechanism (Table 1). Balancing a secure, publicly verifiable and decentralized infrastructure with the need for data confidentiality remains an open challenge, which motivates our research.
| Requirement | Approach | |
| R1 | Public enforcement and transparency of process execution should be guaranteed | The control-flow of the process is managed by smart contracts deployed on public blockchain |
| R2 | The process should be independently auditable with low overhead while guaranteeing authenticity | On-chain information retains publicly available process execution tracking information and hashes with locators for associated encrypted data |
| R3 | Information artifacts should be written in a permanent, tamper-proof and non-repudiable way | Messages are stored in a tamper-proof distributed off-chain file storage, and the resource locators are stored on-chain to keep track of information |
| R4 | Access to information artifacts should be controlled while ensuring their overall integrity and availability | MA-ABE encrypts information artifacts stored in a distributed environment. Decryption is possible only if the requester holds the necessary attributes. |
| R5 | User-defined policies should control access levels for authenticated users with a fine-granular scheme | MA-ABE policies are associated with messages within individual activities |
@@ 3 Background and State of the Art
In this section, we first describe the core concepts underlying our approach, namely blockchain technology and Attribute-Based Encryption (ABE), and then illustrate how these pillars serve as the foundation for existing approaches in the literature relevant to our investigation.
@@ 3.1 Background
A ledger is an append-only singly-linked list of transactions, representing asset transfers among accounts. A blockchain collates ledger segments into blocks, with headers storing block numbers and hashes linking to previous blocks, ensuring transaction order and data integrity. Selected nodes append new blocks and receive cryptocurrency as rewards. The blockchain is replicated across all nodes, ensuring transparency. Consensus protocols and cryptographic primitives enable public verification and decentralization without third-party oversight. Several platforms implement the blockchain protocol (e.g., Ethereum [31]). Some blockchains allow for the deployment of smart contracts [13] which are programs deployed onto the blockchain and executed within a dedicated virtual machine (e.g., the Ethereum Virtual Machine, EVM). Smart contracts functions are invoked through transactions and their core features are code immutability and invocation storage on the blockchain, making such information available for auditing purposes [10]. Smart contracts execution incurs a cost on the invoker. The cost is measured in gas units and depends on the complexity of the called function and the size of the exchanged data. To save on the latter, distributed tamper-proof storage systems are often employed to store the actual data in a non-modifiable file, while a permalink to that file is saved in the smart contract state. The InterPlanetary File System (IPFS) [1] is a typical example of such a system. Based on a peer-to-peer network wherein each node stores chunks of the data, IPFS binds every file with a string used both as a unique identifier and resource locator. The string is based on the hash of the file content itself so that if the file is altered, the permalink does not match. As we will see next, these features fostered the development of blockchain-based PAISs using smart contracts as the backbone for business (process) rule enforcing (in our example, the BPMN choreography is translated into smart contract code, and its steps correspond to invocations to its functions) [7, 9].
@@
ABE is a public key encryption where the ciphertext, an encrypted plaintext, and the corresponding decryption key are linked through custom boolean attributes [25]. In Ciphertext-Policy Attribute-Based Encryption (CP-ABE) [2], a variant of ABE, every user bears a set of attributes encoded in their key.
For example, the owner of the blockchain account 0xB0…1AA1 can be endowed with attributes Patient and PID476948 specifying role and process instance , respectively.
Data owners write policies to be attached to the CP-ABE-encrypted document (the ciphertext). Those policies express conditions for granting in-clear access to the data. Policies are propositional logic formulae employing ABE attributes as literals. The policy is evaluated on attributes associated with a potential reader: Only if the formula is satisfied can the user’s key decrypt the data.
@@
Attributes are typically assigned by one authority: Multi-Authority Attribute-Based Encryption (MA-ABE) [4] is a variant of ABE that removes this single point of failure using a multi-authority method. In MA-ABE, every authority creates a part of the decryption key.
Once collected the user merges all parts to obtain the final key. In this paper, we employ the Ciphertext-Policy variant of MA-ABE [17]. It integrates MA-ABE with the aforementioned policy-based approach.
@@ 3.2 State of the Art
Over the last few years, several solutions that automate collaborative processes using blockchain technology have been designed [27, 9, 18, 29]. Such approaches demonstrated the effectiveness of blockchain-based solutions in improving trust among participants in multi-party collaborations and allowing for monitoring and auditing [10, 8]. These studies enhance the integration of blockchain technology with process management, unlocking security and traceability benefits. However, they primarily focus on the control-flow perspective and lack mechanisms for secure access control to the stored data on public platforms.
A relevant area of research for our work pertains to data privacy and integrity within blockchain systems. Several studies examine the application of encryption techniques to achieve this goal. Hawk [14] employs user-defined private smart contracts to automate the implementation of cryptographic protocols. Pham et al. [24] propose a decentralized storage solution that integrates InterPlanetary File System (IPFS), MA-ABE, and blockchain technology. Hong et al. [12] introduce a data-sharing solution based on decentralized ABE (MA-ABE), blockchain, and IPFS. Bramm et al. [3] present BDABE, an access control mechanism that applies a distributed ABE scheme within a consensus-driven infrastructure for real-world applications. Yan et al. [33] introduce a scheme for fine-grained access control by implementing proxy encryption and decryption while supporting policy hiding and attribute revocation. The encrypted data is stored on IPFS, and the metadata is stored on the blockchain. All these studies, along with the aforementioned [21, 20, 19], cater for secure access control on blockchains but lack the integration of this technology with process management systems.
@@ 4 The CONFETTY Architecture
In this section, we describe the proposed CONFETTY architecture. We first introduce its core components and functionalities (Sect. 4.1). Then, we describe how components interact and manage information to support confidential-preserving blockchain-based process execution while meeting Tables 1, 1, 1, 1 and 1 (Sect. 4.2). Due to space restrictions, we do not enter the technical details of our approach but provide an overarching presentation of our solution’s rationale and key concepts, motivating them with the requirements they aim to meet.
@@ 4.1 Functional Viewpoint
Figure 2
provides an overview of the CONFETTY approach,
@@
(i) The configure functionality helps the process owners specify the process to run on-chain. It requires the registration of process information, including activities to execute, data to exchange, control-flow routing, the expected participants’ roles (like the information enclosed in the choreography diagram in Fig. 1) and how they can operate with confidential data (via MA-ABE policies) to control information access.
(ii) The instantiate functionality lets certifiers authenticate the participants, define their role in the process (e.g., that user 0xB0…1AA1 is a patient), and kickstart a new process instance on-chain by setting its public state.
(iii) The transact functionality allows process participants to interact, following the process, to exchange data at run time. This functionality manages both public and confidential data, updating and enforcing the process state according to the logic in the blockchain.
@@
The functionalities above are realized by the CONFETTY main components depicted in Fig. 3. CONFETTY relies on two external components used as architectural buttresses: (i) A Programmable Blockchain Platform (like Ethereum), (ii) and a Tamper-proof Distributed File Storage (like IPFS). The former maintains the public state, allowing for the creation of logic and constraints while notarizing the executed operations. It comprises two core elements: The Distributed Ledger, storing the transactions, and the Virtual Machine, hosting and running smart contracts. The latter (which we will henceforth name as Distributed FS for brevity) stores business data and thus improves the performance and costs of a blockchain as in common practice. The data is saved on the external storage (upon encryption, if it must remain secret), while a resource locator with the hash of that data is kept on-chain for notarization and future retrieval. Notice that we take inspiration from known patterns adopting Distributed FSs for large data storage [32]. Based on this known pattern, we devise new techniques to preserve secrecy and control access to that process data at runtime. The primary components of CONFETTY are the following: (i) The Process Interface handles all the operations involving the participants and related to the process. This is done by acting as an interface for the blockchain, abstracting from the technical implementation and exposing functionalities to the users. In particular, the Process Interface instantiates the process on the blockchain on given process specification data. To support the instantiate functionality, it provides capabilities for role assignment to participants. (ii) The Process Contract is a smart contract that manages the process instances and their execution. It elaborates and manages all process specifications and their instances’ public states, acting as a software realization both of factory and proxy patterns [32]. It supports the instantiate functionality by enforcing the control flow, the data flow, the assignment of tasks to participants, and the routing of decision points. In this way, the Process Contract updates the public state of the process and generates transparent records that are accessible and verifiable by interested auditors. (iii) The Confidentiality Interface is in charge of handling the participants’ authorizations and the operations over confidential data. It operates as a façade towards the blockchain to store access grants, distributes decryption keys to the users solely based on their roles and involvement in the process, and handles the storage and retrieval of encrypted data. (iv) The Confidentiality Contract is a smart contract handling the notarization of authorizations for confidential data, as well as for their writing and access.
@@
We conclude this subsection with a few considerations about the users and trust model. We assume that the users we mentioned while discussing the functionalities (i.e., process owners, certifiers, process participants, and auditors) possess a blockchain account with a key pair to be able to publicly identify themselves (as with the aforementioned user 0xB0…1AA1) and sign blockchain transactions built through the CONFETTY modules. We also remark that we consider the CONFETTY platform as a trusted party, while users are assumed to be honest but curious. Finally, notice that the involvement of a pre-appointed third-party certifier to authenticate process participants is necessary as a user alone can confirm their identity (e.g., via the aforementioned public-private key scheme) but cannot attest alone to the truthfulness of assertions regarding themselves. The selection and appointment of certifiers transcends the goals of this paper, but in our motivating scenario, e.g., certifiers can be trusted third parties like the Public Ministry of Health (for hospital personnel), the Public Registry (for citizens), and the Insurance Registrar (for agencies), or subsidiaries thereof. In our architecture, certifiers operate like blockchain oracles [23].